XYZ company operates in a perfectly competitive market where the current market price is $10. Currently the firm is producing 200 units at an average variable cost of $8, an average total cost of $12 and a marginal cost of $10.

a. Is the situation described above a short-run equilibrium for XYZ? Explain.

b. What is XYZ’s profit/loss?

c. What is XYZ’s producer surplus?

d. Should XYZ continue to produce in this situation? Explain.

e. Assuming that XYZ is ‘average’ (i.e., neither more nor less efficient than other firms in the market), what would you expect to happen to market price in the long run? Explain.
Answer #1

a) Total Variable Cost = AVC * No. of Units of Output = $8 * 200 = $1600

Total Cost = ATC * No. of Units of Output = $12 * 200 = $ 2400

Fixed Cost = Total Cost - Total Variable Cost = $2400 - $1600 = $800

Since firm is incurring Fixed Costs, this means it is operating in short run. Because in long run all costs are variable and there is no fixed cost.

b) Profit / Loss = Total Revenue - Total Cost

Total Revenue = Price * Quantity = $10 * 200 = $2000

Total Cost = $2400

Loss = $2000 - $2400 = -$400

c) In the short run, a perfectly competitive firm's producer surplus is equal to Total Revenue minus Variable Cost. Total Revenue = $2000. Total Variable Cost = $1600. Producer Surplus = $2000 - $1600 = $400.

d) Yes, XYZ should continue to produce. A firm should continue to produce in the short run until P > AVC. Here, Price per unit is $10 and Average Variable Cost is $8. Since Price is greater than AVC, so XYZ should continue to produce.